# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

### v3

@@ IV-B Dynamic Rank Allocation
The fixed rank configuration assigns uniform ranks to modules across different layers and components, resulting in communication inefficiency due to its inability to adapt to the varying contributions of layers and components to model performance. This inefficiency manifests in two aspects: the fixed rank allocation yields suboptimal performance under similar communication overhead and requires higher communication costs to achieve comparable performance with the dynamic rank allocation.
@@
IV-B1 Generation of Local Rank Masks
@@
Mag 63.89 80.24 64.32 76.89 1$\times$
Grad 63.31 79.59 63.68 76.05 0.97$\times$
Mixed 63.79 79.87 63.94 76.42 1.11$\times$
Sensitivity 63.89 80.06 64.21 76.70 1.30$\times$
@@
IV-B2 Arbitration of Global Rank Masks
here, ${K}$ denotes the set of participating clients. The arbitrated global rank masks are broadcast to all clients and used to optimize communication parameters between the server and clients. As shown in Table II, we compare two arbitration strategies: FedARA collects local rank masks from clients and performs arbitration on the server; and FedARA-global generates global rank masks directly from the aggregated model. Although the communication overhead is similar for both strategies, FedARA achieves better performance on the 20News dataset [25], highlighting the effectiveness of leveraging local rank information for centralized arbitration.
IV-B3 Communication with Arbitrated Masks
@@ IV-C Rank-based Module Pruning
Although the dynamic parameter allocation method with truncated SVD modules reduces communication overhead in FL by limiting the upload and download of less critical ranks, it does not decrease the number of trainable parameters during local training. This limitation arises because frameworks such as PyTorch and Transformers manage trainable parameters at the matrix level, preventing selective updates at the rank level. Fine-grained sparse gradient updates require additional computation overhead. To address this, we introduce a rank-based module pruning mechanism that safely removes inactive SVD modules in a structured manner.
@@ V Experimental Setups
FedARA extends AdaLoRA [28], originally designed for dynamic parameter configuration in centralized training, to the federated learning setting. Built on Hugging Face’s Transformers library, it supports a broad range of pre-trained NLP models. FedARA adopts a random client selection mechanism, which can be substituted with more advanced strategies [29, 30]. Communication between clients and the server is synchronized via weighted averaging [7], maintaining compatibility with diverse optimization and aggregation methods [31, 32, 33, 34, 35, 36]. Moreover, FedARA incorporates model quantization techniques such as QLoRA [37] and QA-LoRA [38] to reduce computational and memory overhead, potentially lowering communication costs.
Hardware. Following prior FL research [23], all experiments are conducted via emulation on a laptop equipped with an RTX 4070 GPU. Local training time is measured on three edge devices—AGX Orin, Orin Nano, and Raspberry Pi 5—chosen for their representativeness of common resource-constrained hardware. These results are integrated into our emulation framework to estimate total training time in a homogeneous FL setting. In line with previous studies [23], we assume a communication bandwidth of 1MB/s between the server and clients, reflecting realistic conditions across IoT, home WiFi, and cellular networks.
TABLE III: Different dataset settings and sizes in sequence classification and sumarization tasks of NLP.
@@
Datasets and Division. FedARA is evaluated on sequence classification and summarization tasks using five diverse NLP datasets, as summarized in Table III. Classification tasks include: (1) 20News [25], comprising documents from 20 different newsgroups; (2) Semeval [40], covering various NLP and semantic evaluation tasks; (3) AG News [41], containing news articles from four categories; (4) News Category [27], a dataset of diverse news headlines, where we focus on the top 15 most frequent labels, covering approximately 70% of the data. Summarization task: (5) CNN/DailyMail [42], a widely used dataset for abstractive summarization. Each classification dataset is split into training, validation, and test sets in an 8:1:1 ratio. We fine-tune a DistilBERT model on the News Category dataset (top-15 labels) and use the resulting model to generate pseudo-labels for the summarization dataset. To simulate non-IID conditions in a federated setup, we adopt Dirichlet-based partitioning with $\alpha\in\{1,0.1,0.01\}$. Following FedAvg [7], we also consider a pathological non-IID setting where each client holds samples from only 1–2 labels. For IID settings, we set $\alpha=1000$ to approximate uniform label distribution across clients.
Model. We adopt three widely used transformer-based models for NLP tasks. For classification, we use BERT-base [3], which consists of 12 transformer layers, and its compressed variant DistilBERT-base [2], which retains most of BERT’s performance with only 6 layers via knowledge distillation. For summarization, we employ BART-base [4], an encoder–decoder model designed for sequence-to-sequence generation. All selected models are based on the transformer architecture, which remains foundational for even the latest NLP models [10]. Our work focuses on federated adaptive rank allocation for transformer models. While this study evaluates BERT, DistilBERT, and BART across classification and summarization tasks to ensure generality and feasibility on resource-constrained devices, our method is readily applicable to newer architectures and broader NLP tasks.
TABLE IV: Comparison of FedPEFT methods across different models and datasets. In A (B), A represents the accuracy under non-IID data (Pathological distribution), while B denotes the relative accuracy degradation compared to IID data.
@@
  • FedAdapter-p: inserts adapter modules only in the feed-forward layers [12]. We exclude AdaFL [23] as it is a IIIparameter-reduced version of FedAdapter-p without targeted non-IID optimization.
  • SLoRA: employs a two-stage training scheme [19], where the first stage applies sparse fine-tuning (SFT) to improve initialization for LoRA in the second stage. We allocate 10% of the total FL rounds to the SFT stage.
@@
  • FFA-LoRA: updates only the $B$ matrix in LoRA while freezing $A$ [26, 22, 43]. To mitigate reduced expressiveness, we also evaluate FFA-LoRA-dr, which employs orthogonal initialization for $A$ and doubles the rank.
@@ VI-A Performance under non-IID data
FedARA effectively mitigates performance degradation under non-IID settings. FedARA consistently outperforms existing methods across four datasets and two models. Using DistilBERT, FedARA achieves an average accuracy improvement of 8.49% over FedLoRA and 11.58% over FedAdapter-h, while also surpassing FFA-LoRA-dr and FFA-LoRA by 6.95% and 7.75%, respectively. The largest gain 37.59% is observed on the Semeval dataset, which has more labels and fewer samples, whereas the smallest gain 1.47% occurs on AG News, with fewer labels and more samples. Moreover, FedARA significantly outperforms SLoRA and FeDeRA, which rely on modified LoRA initialization strategies.
We further conduct experiments under varying degrees of non-IID data distributions by adjusting the Dirichlet parameter $\alpha$, where smaller $\alpha$ indicates higher non-IIDness. The results are illustrated in Figure 7. As $\alpha$ decreases, the performance of FedPEFT shows a clear downward trend, which is consistent with the observations reported in [19, 20]. Moreover, FedARA consistently outperforms other FedPEFT methods across different $\alpha$ values, with the performance gap becoming particularly significant under highly skewed non-IID settings.
@@
FedARA significantly improves communication efficiency. FedARA reduces the communication overhead while maintaining superior performance. As shown in Figure 8, FedARA begins with a similar initial communication cost (r=12) as baseline methods on the BERT model and AG News dataset, but quickly drops it to 131.42 MB—71.17% lower than its initial warm-up cost. SLoRA and FeDeRA incur the same communication cost as FedLoRA. Table IV further shows that even when FedARA’s initial rank is twice that of FedLoRA, its total communication overhead remains nearly 20% lower than that of FedLoRA and FFA-LoRA-dr. When initial ranks are equal, FedARA reduces communication overhead by approximately 2.40$\times$. FFA-LoRA also incurs about 20% more overhead than FedARA. Overall, FedARA achieves the lowest communication cost across settings while consistently outperforming all baselines.
@@
Figure 9 shows the final rank distribution of FedARA on the DistilBERT model and 20News dataset. Starting from a uniform rank of 12, FedARA adaptively reduces the average rank to 3, effectively balancing performance and communication cost. Darker blocks in Figure 8 represent higher ranks, mainly in deeper layers and key components like query, key, f1, and f2, highlighting their role in preserving important model information—consistent with our earlier findings (Figure 2c).
@@ VI-B Training Time on Advanced Edge Devices
We evaluate the training efficiency of FedARA and baselines on various edge devices. On Raspberry Pi 5 (RPi 5), local training time per batch (batch size = 4) is 1.00 s for DistilBERT and 2.01 s for BERT. AGX Orin and Orin Nano achieve significantly faster training under DistilBERT, with 6.67$\times$ and 5.56$\times$ speedups over RPi 5, respectively; for BERT, the speedups increase to 8.74$\times$ and 6.70$\times$.
@@
Figure 10 compares total training time for the DistilBERT model across devices. On RPi 5, FedARA reduces training time to 68 minutes, representing 27.65% and 24.38% reductions over FedLoRA and FFA-LoRA-dr, respectively. On Orin Nano and AGX Orin, FedARA further lowers training time to 25 and 23 minutes, achieving up to 48.90% reduction over FedLoRA and over 12% compared to FFA-LoRA. These results suggest that limited compute on edge devices may diminish the benefits of reduced communication overhead. This observation aligns with Figure 2d, where high-end devices are communication-bound, while low-end devices remain compute-bound. We exclude comparisons with SLoRA and FeDeRA due to their noticeably higher training and computational costs compared to FedLoRA. Specifically, SLoRA incurs overhead comparable to full fine-tuning when using optimizer masking approach [18], while FeDeRA introduces additional model-size-dependent costs through SVD.
@@ VI-C Ablation Experiments
Ablation study of dynamic rank allocation. To examine the role of dynamic rank allocation, we evaluate FedARA-r8 and FedARA-r4, which have initial ranks of 8 and 4, respectively. As shown in Figure 11, FedARA-r8 improves average performance by 0.45% and reduces total communication overhead by nearly 20% compared to FedSVD. Meanwhile, FedARA-r4 achieves comparable or superior performance while reducing communication overhead by approximately 2.20$\times$. These results confirm that dynamic rank allocation maximizes communication efficiency.
Additionally, we assess the per-round communication overhead across FedLoRA, FedSVD, and FedARA. As illustrated in Figure 12, all methods begin with a similar overhead of around 75.98 MB. While FedLoRA and FedSVD maintain constant communication costs, FedARA’s dynamic rank allocation progressively reduces costs, stabilizing at 22.17 MB after 50 rounds—a 70.81% and 70.82% reduction compared to FedLoRA and FedSVD. This suggests that FedARA’s fine-tuned model requires fewer trainable parameters, offering potential benefits for minimizing storage costs of PEFT modules for various downstream tasks.
Ablation study of rank-based module pruning. Rank-based module pruning does not impact performance, as it preserves the same parameters for communication and global aggregation. We evaluate its effect on local training time and GPU usage on Semeval and News Category datasets. As illustrated in Figure 13 and 14, while FedARA initially exhibits a slight increase in local training time, it significantly reduces over time. Compared to FedLoRA, FedARA lowers average local training time by 10.81% and reduces GPU memory consumption by 31.67% per round, demonstrating its efficiency in minimizing computational and storage overhead.
@@ VI-E Local Resource Utilization
Memory Footprint. Although SVD adaptation introduces an additional diagonal matrix, the memory overhead remains negligible due to the small rank $r$ relative to the hidden layer size. As shown in Figure 16, integrating SVD modules and dynamic rank allocation in FedARA leads to only a slight increase in peak memory usage compared to FedLoRA—2.42% for DistilBERT and 4.40% for BERT on Orin Nano. Moreover, FedARA progressively reduces memory usage per round during federated fine-tuning, confirming its suitability for memory-constrained devices.
Energy Consumption. Figure 17 shows the energy consumption on Orin Nano (15W) over 100 rounds with 10 clients per round. FedARA achieves the lowest energy usage—approximately $2.3\times 10^{5}$ J for DistilBERT and $4.4\times 10^{5}$ J for BERT—representing a 46.95% reduction compared to FedLoRA and around 9% compared to FFA-LoRA. These results demonstrate FedARA’s ability to reduce energy consumption, extending device runtime and lowering deployment costs for sustainable use in resource-constrained environments.
@@ VII Related Work
Non-IID Data. Non-IID data in FL hinders model aggregation and global generalization. Traditional methods—such as client selection, optimization, and aggregation—have been effective in mitigating these effects [29, 30, 31, 32, 33, 34, 35, 36]. Recent work extends this line to pre-trained models within the FedPEFT framework. SLoRA [19] adopts a two-stage strategy: sparse fine-tuning for initialization, followed by LoRA adaptation. FeDeRA [20] enhances LoRA initialization via SVD. Beyond initialization, recent studies explore asymmetric training by updating only matrix $B$ while freezing matrix $A$ to reduce non-IID impact [43, 26, 22]. FedARA further addresses non-IID challenges by modifying both the structure and initialization via truncated SVD adaptation, effectively slowing client drift.
@@
Dynamic Parameter Configurations. Fixed rank configurations lead to dual losses of suboptimal performance and communication inefficiency. In centralized training, AdaLoRA [28] introduces a sensitivity-based dynamic rank allocation method. In the context of FedPEFT, AdaFL [23] expands parameters via sideline trials on other clients, but introduces inference latency and additional communication costs. SA-FedLoRA [21] employs a two-stage simulated annealing approach: full fine-tuning with regularization to control model drift, followed by adaptive LoRA rank adjustment. However, the high cost of full fine-tuning limits its practicality on resource-constrained devices. FedARA replaces full fine-tuning with truncated SVD adaptation for resource efficiency, and leverages rank masks to reduce communication without extra clients, enabling a more scalable solution.

### v4

@@ IV-B Dynamic Rank Allocation
The fixed-rank configuration assigns uniform ranks to modules across different layers and components, resulting in communication inefficiency due to its inability to adapt to the varying contributions of layers and components to model performance. This inefficiency manifests in two aspects: the fixed rank allocation yields suboptimal performance under similar communication overhead and requires higher communication costs to achieve comparable performance with the dynamic rank allocation.
@@
IV-B1 Generation of Local Rank Masks (MaskGen)
@@
[…]
As illustrated in Figure 6, each SVD module initially contains $r$ triplets, which can be inserted into six different components. The local rank mask of each client is generated based on the sorting of all triplets. Specifically, if the importance of a rank falls within the $\textit{top-}b^{(t)}$, it is marked as True in the mask; otherwise, it is marked as False. Each client uploads these masks to the server along with the parameters.
IV-B2 Arbitration of Global Rank Masks (FedArb)
here, K{K} denotes the set of participating clients. The arbitrated global rank masks are broadcast to all clients and used to optimize communication parameters between the server and clients. As shown in Table II, we compare two arbitration strategies: FedARA collects local rank masks from clients and performs arbitration on the server; and FedARA-global generates global rank masks directly from the aggregated model. Although the communication overhead is similar for both strategies, FedARA achieves better performance on the 20News dataset [18], highlighting the effectiveness of leveraging local rank information for centralized arbitration.
@@ IV-B2 Arbitration of Global Rank Masks (FedArb)
IV-B3 Communication with Arbitrated Masks (CommPru)
@@ IV-C Rank-based Module Pruning
Although the dynamic parameter allocation method with truncated SVD modules reduces communication overhead in FL by limiting the upload and download of less critical ranks, it does not decrease the number of trainable parameters during local training. This limitation arises because frameworks such as PyTorch and Transformers manage trainable parameters at the matrix level, preventing selective updates at the rank level. Fine-grained sparse gradient updates require additional computation overhead. To address this, we introduce a rank-based module pruning mechanism (RankDet in Algorithm 1) that safely removes inactive SVD modules in a structured manner.
@@ V Experimental Setups
FedARA extends AdaLoRA [42], originally designed for dynamic parameter configuration in centralized training, to the federated setting. Built on Hugging Face’s Transformers library, it supports a broad range of pre-trained NLP models. FedARA adopts a random client selection mechanism, which can be substituted with more advanced strategies [45, 10]. Communication between clients and the server is synchronized via weighted averaging [25], maintaining compatibility with diverse optimization and aggregation methods [21, 16, 26, 39, 29]. Moreover, FedARA incorporates model quantization techniques such as QLoRA [6] and QA-LoRA [36] to reduce computational and memory overhead, potentially lowering communication costs.
TABLE III: Different dataset settings and sizes in sequence classification and summarization tasks of NLP.
@@
Hardware. Following prior FL research [4], all experiments are conducted via emulation on a laptop equipped with an RTX 4070 GPU. Local training time is measured on three edge devices, AGX Orin, Orin Nano, and Raspberry Pi 5, chosen for their representativeness of common resource-constrained hardware. These results are integrated into our emulation framework to estimate total training time in a homogeneous FL setting. In line with previous studies [4], we assume a communication bandwidth of 1MB/s between the server and clients, reflecting realistic conditions across IoT, home WiFi, and cellular networks.
Datasets and Partitioning. FedARA is evaluated on sequence classification and summarization tasks using five diverse NLP datasets, as summarized in Table III. Classification tasks include: (1) 20News [18], comprising documents from 20 different newsgroups; (2) Semeval [11], covering various NLP and semantic evaluation tasks; (3) AG News [43], containing news articles from four categories; (4) News Category [27], a dataset of diverse news headlines, where we focus on the top 15 most frequent labels, covering approximately 70% of the data. Summarization task: (5) CNN/DailyMail [12], a widely used dataset for abstractive summarization. Each classification dataset is split into training, validation, and test sets in an 8:1:1 ratio. We fine-tune a DistilBERT model on the News Category dataset (top-15 labels) and use the resulting model to generate pseudo-labels for the summarization dataset. To simulate non-IID conditions in a federated setup, we adopt Dirichlet-based partitioning with $\alpha\in\{1,0.1,0.01\}$. Following FedAvg [25], we also consider a pathological non-IID setting where each client holds samples from only 1-2 labels. For IID settings, we set $\alpha=1000$ to approximate a uniform label distribution across clients.
Model. We adopt three widely used transformer-based models for NLP tasks. For classification, we use BERT-base [7], which consists of 12 transformer layers, and its compressed variant DistilBERT-base [30], which retains most of BERT’s performance with only 6 layers via knowledge distillation. For summarization, we employ BART-base [20], an encoder-decoder model designed for sequence-to-sequence generation. All selected models are based on the transformer architecture, which remains foundational for even the latest NLP models [8]. Our work focuses on federated adaptive rank allocation for transformer models. While this study evaluates BERT, DistilBERT, and BART across classification and summarization tasks to ensure generality and feasibility on resource-constrained devices, our method is readily applicable to newer architectures and broader NLP tasks.
TABLE IV: Comparison of FedPEFT methods across different models and datasets. In N1 (N2), N1 represents the accuracy under non-IID data (Pathological distribution), while N2 denotes the relative accuracy degradation compared to IID data.
@@
  • FedAdapter-p: inserts adapter modules only in the feed-forward layers [28]. We exclude AdaFL [4] as it is a parameter-efficient variant of FedAdapter-p and does not explicitly address non-IID data heterogeneity.
  • SLoRA: employs a two-stage training scheme [1], where the first stage applies sparse fine-tuning to improve initialization for LoRA in the second stage. We allocate 10% of the total FL rounds to the sparse stage.
@@
  • FFA-LoRA: updates only the $B$ matrix in LoRA while freezing AA [41, 31, 44]. To mitigate reduced expressiveness, we also evaluate FFA-LoRA-dr, which employs orthogonal initialization for AA and doubles the rank.
@@ VI-A Performance under non-IID data
FedARA effectively mitigates performance degradation under non-IID settings. FedARA consistently outperforms existing methods across four datasets and two models. Using DistilBERT, FedARA achieves an average accuracy improvement of 8.49% over FedLoRA and 11.58% over FedAdapter-h, while also surpassing FFA-LoRA-dr and FFA-LoRA by 6.95% and 7.75%, respectively. The largest gain of 37.59% is observed on the Semeval dataset, which has more labels and fewer samples, whereas the smallest gain of 1.47% occurs on AG News, with fewer labels and more samples. Moreover, FedARA significantly outperforms SLoRA and FeDeRA, which rely on modified LoRA initialization strategies.
We further conduct experiments under varying degrees of non-IID data distributions by adjusting the Dirichlet parameter $\alpha$, where a smaller $\alpha$ indicates higher non-IIDness. The results are illustrated in Figure 7. As $\alpha$ decreases, the performance of FedPEFT shows a clear downward trend, which is consistent with the observations reported in [1, 37]. Moreover, FedARA consistently outperforms other FedPEFT methods across different $\alpha$ values, with the performance gap becoming particularly significant under highly skewed non-IID settings.
@@
[…]
Figure 9 shows the final rank distribution of FedARA on the DistilBERT model and 20News dataset. Starting from a uniform rank of 12, FedARA adaptively reduces the average rank to 3, effectively balancing performance and communication cost. Darker blocks in Figure 8 represent higher ranks, mainly in deeper layers and key components like query, key, f1, and f2, highlighting their role in preserving important model information, consistent with our earlier results in Figure 2c.
@@ VI-B Training Time on Advanced Edge Devices
[…]
@@
Figure 10 compares total training time for the DistilBERT model across devices. On RPi 5, FedARA reduces training time to 68 minutes, representing 27.65% and 24.38% reductions over FedLoRA and FFA-LoRA-dr, respectively. On Orin Nano and AGX Orin, FedARA further lowers training time to 25 and 23 minutes, achieving up to 48.90% reduction over FedLoRA and over 12% compared to FFA-LoRA. These results suggest that limited computing resources on edge devices may diminish the benefits of reduced communication overhead. This observation aligns with Figure 2d, where high-end devices are communication-bound, while low-end devices remain compute-bound. We exclude comparisons with SLoRA and FeDeRA due to their noticeably higher training and computational costs compared to FedLoRA. Specifically, SLoRA incurs overhead comparable to full fine-tuning when using an optimizer masking approach [2], while FeDeRA introduces additional model-size-dependent costs through SVD.
@@ VI-C Ablation Experiments
[…]
@@
Additionally, we assess the per-round communication overhead across FedLoRA, FedSVD, and FedARA. As illustrated in Figure 12, all methods begin with a similar overhead of around 75.98 MB. While FedLoRA and FedSVD maintain constant communication costs, FedARA’s dynamic rank allocation progressively reduces costs, stabilizing at 22.17 MB after 50 rounds, a 70.81% and 70.82% reduction compared to FedLoRA and FedSVD. This suggests that FedARA’s fine-tuned model requires fewer trainable parameters, offering potential benefits for minimizing storage costs of PEFT modules for various downstream tasks.
Ablation study of rank-based module pruning. Rank-based module pruning does not impact performance, as it preserves the same parameters for communication and global aggregation. We evaluate its effect on local training time and GPU usage on Semeval and News Category datasets. As illustrated in Figure 13 and 14, while FedARA initially exhibits a slight increase in local training time, it significantly reduces local training time over time. Compared to FedLoRA, FedARA lowers average local training time by 10.81% and reduces GPU memory consumption by 31.67% per round, demonstrating its efficiency in minimizing computational and storage overhead.
@@ VI-E Local Resource Utilization
Memory Footprint. Although SVD adaptation introduces an additional diagonal matrix, the memory overhead remains negligible due to the small rank $r$ relative to the hidden layer size. As shown in Figure 16, integrating SVD modules and dynamic rank allocation in FedARA leads to only a slight increase in peak memory usage compared to FedLoRA, 2.42% for DistilBERT and 4.40% for BERT on Orin Nano. Moreover, FedARA progressively reduces memory usage per round during federated fine-tuning, confirming its suitability for memory-constrained devices.
Energy Consumption. Figure 17 shows the energy consumption on Orin Nano (15W) over 100 rounds with 10 clients per round. FedARA achieves the lowest energy usage, approximately $2.3\times 10^{5}$ J for DistilBERT and $4.4\times 10^{5}$ J for BERT, representing a 46.95% reduction compared to FedLoRA and around 9% compared to FFA-LoRA. These results demonstrate FedARA’s ability to reduce energy consumption, thereby extending device runtime and lowering deployment costs for sustainable operation in resource-constrained environments.
@@ VII Related Work
Non-IID Data. Non-IID data in FL hinders model aggregation and global generalization. Traditional methods, such as client selection, optimization, and aggregation, have been effective in mitigating these effects [45, 10, 21, 16, 26, 39, 29]. Recent work extends this line to pre-trained models within the FedPEFT framework. SLoRA [1] adopts a two-stage strategy: sparse fine-tuning for initialization, followed by LoRA adaptation. FeDeRA [37] enhances LoRA initialization via SVD. Beyond initialization, recent studies explore asymmetric training by updating only matrix $B$ while freezing matrix $A$ to reduce non-IID impact [44, 41, 31]. FedARA further addresses non-IID challenges by modifying both the structure and initialization via truncated SVD adaptation, effectively slowing client drift.
@@
Dynamic Parameter Configurations. Fixed-rank configurations lead to dual losses of suboptimal performance and communication inefficiency. In centralized training, AdaLoRA [42] introduces a sensitivity-based dynamic rank allocation method. In the context of FedPEFT, AdaFL [4] expands parameters via sideline trials on other clients, but introduces inference latency and additional communication costs. SA-FedLoRA [38] employs a two-stage simulated annealing approach: full fine-tuning with regularization to control model drift, followed by adaptive LoRA rank adjustment. However, the high cost of full fine-tuning limits its practicality on resource-constrained devices. FedARA replaces full fine-tuning with truncated SVD adaptation for resource efficiency, and leverages rank masks to reduce communication without extra clients, enabling a more scalable solution.